OpenStudy (jessiegonzales):

MEDAL!?!?!?! 1. DF bisects

OpenStudy (jessiegonzales):

@zepdrix

OpenStudy (kagıtucak):

Since FG is a angle bisector and FE and FG are perpendicular at their edges n+t must equal to 2n-6. (always valid rule) n+5=2n-6 11=n

OpenStudy (foxracer62):

n+5=2n-6 11=n

OpenStudy (foxracer62):

2(11)-6=FG 22-6=FG 16=FG